OpenStudy (anonymous):

a line through (1,-4) and (-5,2) is perpendicular to the line through (8, y) and (-6, -7) Find y

OpenStudy (anonymous):

find the slopes from these points and relate them in this manner m1 = -1/m2 you will get y

OpenStudy (anonymous):

(1,-4) and (-5,2) can you find the slope from these points???

OpenStudy (anonymous):

-6/4?? correct??

hartnn (hartnn):

numerator is correct, try again for denominator....

hartnn (hartnn):

(-4-2) /(1-(-5)) =... ?

OpenStudy (anonymous):

6 also??

hartnn (hartnn):

yes, so the slope is -6/6 =-1 then what will be the slope of perpendicular line ??

hartnn (hartnn):

m1 = -1/m2 m2=-1

OpenStudy (jhannybean):

A line that is perpendicular to another line has the form \[\large m= -\frac1m\] original line (m1) \[m_{1} = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}=\frac{2+4}{-5-1}=\frac{6}{-6}=-1\] andif the slope of the perpendicular line is the negative opposite reciprocal, your outcome would have to be the opposite of the slope we found, \[\large m_{2}=\frac{-7-y}{-6-8}= \frac{?}{-14}\] what number would you subtract from 7 to give you a positive outcome at the end?